OpenStudy (anonymous):

8y - 3 = 5(y - 1) + 3y

OpenStudy (nerdgamer_):

All you really need to do is combine like terms

OpenStudy (nerdgamer_):

But this is a sorta different question so I dont know if I can help so Okay

OpenStudy (anonymous):

11y-3=5(1) im kind of lost ..

OpenStudy (pratyush5):

8y - 3 = 5(y - 1) + 3y 8y - 3 = 5y - 5 + 3y 8y - 3 = 8y - 5 -3 = -5 This is never possible, so there is no value of y that satisfies this. No solution